Look up the wavelength of X-ray commonly used in medical diagnostic X-rays and in X-rays emitted during a supernova. From this wavelength, calculate the energy of these X-rays. In addition, calculate the frequency of the X-rays.
1
Expert's answer
2018-05-08T10:55:08-0400
1) X-ray commonly used in medical diagnostic X-rays
λ=〖10〗^(-11) m.
E=hc/λ=(6.63∙〖10〗^(-34) (3∙〖10〗^8))/〖10〗^(-11) =2∙〖10〗^(-14) J
f=c/λ=((3∙〖10〗^8))/〖10〗^(-11) =3∙〖10〗^19 Hz
2) X-rays emitted during a supernova
λ=〖10〗^(-8) m.
E=hc/λ=(6.63∙〖10〗^(-34) (3∙〖10〗^8))/〖10〗^(-8) =2∙〖10〗^(-17) J
f=c/λ=((3∙〖10〗^8))/〖10〗^(-8) =3∙〖10〗^16 Hz